Exercise 8-12 a-b Elburn Supply Co. has the following transactions related to notes receivable during the last 2 months of 2020. The company does not make entries to accrue interest except at December 31. Nov. 1 Loaned $30,000 cash to Manny Lopez on a 12 month, 10% note. Dec. 11 Sold goods to Ralph Kremer, Inc., receiving a $6,750, 90-day, 8% note. 16 Received a $4,000, 180 day, 9% note in exchange for Joe Fernetti’s outstanding accounts receivable. 31 Accrued interest revenue on all notes receivable. Record the collection of the Lopez note at its maturity in 2021. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

I am having trouble with the record the collection of the Lopez note at its maturity in 2021
Answer #1
Date Account title Debit credit
31 Oct 2021 cash 33000
Interest receivable 500
Interest revenue 2500
Note receivable 30000

**Interest accrued on Lopez note for a 12 month period = 30000*10%= 3000. out of which 2month interest was accrued in 2020 [1Nov-31Dec] amounting to 30000*10%*2/12 = 500 so rest is accrue in 2021 which is equals to 3000-500 = 2500